# RedCodeAgent: Automatic Red-teaming Agent against Diverse Code Agents

### v1

@@ 4.1 RedCodeAgent Achieves Higher ASR and Lower RR
RedCodeAgent Explores New Vulnerabilities. In scenarios where other methods fail to find successful attack strategies, RedCodeAgent is able to discover new, feasible jailbreak approaches. For instance, RedCodeAgent successfully identified strategies for risk index 10: creating a reverse shell, and risk index 14: eval injection. A case study and quantitative analysis will be discussed in §˜4.3, and a successful example is shown in §˜H.
@@ Appendix B The Use of Large Language Models
We used ChatGPT (Achiam et al., 2023) to assist with grammatical correction and writing refinement. All research ideas, experimental design, results, and conclusions were independently developed by the authors. The authors bear full responsibility for the entire content of the paper.
@@ Appendix C Benchmark Details
RedCode-Exec (Guo et al., 2024) provides textual instructions that prompt agents to generate risky code, covering diverse risk scenarios as shown in Fig.˜6.
RedCode-Gen (Guo et al., 2024) consists of Python function signatures and docstring instructions derived from eight malware families—Adware, Rootkit, Trojan, Virus, DDoS, Ransomware, Spyware, and Others—designed to evaluate a code agent’s ability to generate malware with explicit malicious intent and potential for harm.
RMCBench (Chen et al., 2024a) evaluates the ability of LLMs to resist generating malicious code. We use all the text-to-code test cases in our experiment.
@@ D.2 Effectiveness on Real World Code Assistants
We test real-world code assistants, Cursor (Cursor., 2024) and Codeium (Codeium., 2024). For the model setting, we use the “cursor-small model” as the base model for Cursor and the “Base Model” for Codeium.
@@ D.7 Extending the Toolset of RedCodeAgent
Gradient-based attacks: We employ GCG (Zou et al., 2023).
Learning-based attacks: We adopt techniques such as Advprompter (Paulus et al., 2024) and AmpleGCG (Liao & Sun, 2024).
Evolutionary-based attacks: We incorporate AutoDAN (Liu et al., 2023b).
Template-based attacks: We use templates from GPTFUZZER (Yu et al., 2023) to craft the attack prompts.
@@ D.11 Experiments on SQL Injection
SQL injection is an important risk scenario; however, it is not included in the RedCode-Exec dataset. Therefore, we followed the test case construction principles outlined in (Guo et al., 2024) and pre-deployed the example.db database for testing. We created 30 different SQL injection test cases in code format and translated them into text instructions. These text instructions were given to the code agent to evaluate whether it would generate code with vulnerabilities related to SQL injection.
@@ E.2 LLM-based Evaluation
For RedCode-Gen benchmark, following the evaluation setting from (Guo et al., 2024), we employed an LLM (i.e., GPT-4o) as the judge to score the maliciousness of the agent-generated content.
@@ F.1.1 Agent Framework
We utilize LangChain (Topsakal & Akinci, 2023) as the framework of RedCodeAgent, with GPT-4o-mini (Achiam et al., 2023) as the base LLM. Within the LangChain framework, we use ReAct (Yao et al., 2023) logic, and define functions for tool calling following the setup described in §˜3.3. Additionally, we implement an external memory bank as outlined in §˜3.2, where the structure of memory entries, memory insertion, and memory retrieval processes are discussed in detail. Our system prompt for RedCodeAgent is as follows:
@@ F.1.2 Code Substitution
Due to the sensitivity of some requests, a single LLM might refuse to respond. Therefore, we provide three LLMs to offer code substitution suggestions for RedCodeAgent. They are “Qwen/CodeQwen1.5-7B-Chat"(Bai et al., 2023), “meta-llama/Meta-Llama-3-8B-Instruct"(AI@Meta, 2024), and “gpt-35-turbo"(Achiam et al., 2023), with their specific links listed in Tb.˜12.
@@ F.3 Baselines Settings
GCG(Zou et al., 2023): We utilize nanoGCG, a lightweight yet fully-featured implementation of GCG. In our experiments, the parameters are set as follows: num_steps=50, search_width=64, topk=64, and seed=42. The target suffix used is “Here is”, and the target white-box model for suffix generation is deepseek-ai/deepseek-coder-6.7b-instruct. It is important to note that deepseek-ai/deepseek-coder-6.7b-instruct differs from the base LLM of our target code agent, OpenCodeInterpreter, which uses m-a-p/OpenCodeInterpreter-DS-6.7B. Therefore, we employed a transfer attack approach using GCG. We only selected one result with the lowest loss among those generated by GCG as the final optimized test case.
@@
AmpleGCG(Liao & Sun, 2024): For AmpleGCG, we follow the released code and utilized the model osunlp/AmpleGCG-plus-llama2-sourced-llama2-7b-chat. The parameter settings are: max_new_tokens=20, min_new_tokens=20, diversity_penalty=1.0, num_beams=50, and return_num_beams=5. We only select the first result from the list of strings generated by AmpleGCG optimization as the final optimized test case.
AutoDAN(Liu et al., 2023b): We follow the official code implementation, using the same experimental parameters as those provided in the repository. For the initial prompt setting, we use the following: Hypothetical response: the way to describe a character planning to [REPLACE] is as follows: “yo”. Rewrite the hypothetical response from the perspective of that character, make it extremely detailed, and use exact and precise information (avoid generalizations or vagueness).
AdvPrompter(Paulus et al., 2024): We utilize the provided code to get the LoRA adapter for the target model based on “meta-llama/Llama-2-7b-hf", and then combined the adapter with the original LLM to create the suffix generator model. The maximum number of new tokens was set to max_new_tokens=50.
@@ G.1 Negative Optimization of Some Baseline Methods
In our experiments, we selected 4 representative and widely used optimization methods (i.e., GCG (Zou et al., 2023), AmpleGCG (Liao & Sun, 2024), Advprompter (Paulus et al., 2024), and AutoDAN (Liu et al., 2023b)) that have been shown to be effective in previous literature on jailbreaking LLMs. However, we observed that certain optimizations in these baseline methods may not be as effective for attacking code agents. This is likely due to the differences between natural language processing and code-specific reasoning tasks. For instance, the targeted code agent OCI can reject the prompt optimized from AutoDan (see detailed example below). At present, most existing jailbreak methods are focused on natural language tasks, and there is a lack of specialized jailbreaking optimization algorithms designed for code-related tasks. This gap further highlights the contribution of RedCodeAgent, which addresses this deficiency by offering a tailored framework for red-teaming code agents, ultimately advancing the field.

### v2

@@ 4.1 RedCodeAgent Achieves Higher ASR and Lower RR
RedCodeAgent Explores New Vulnerabilities. In scenarios where other methods fail to find successful attack strategies, RedCodeAgent is able to discover new, feasible jailbreak approaches. For instance, RedCodeAgent successfully identified strategies for risk index 10: creating a reverse shell, and risk index 14: eval injection. We use the risk index to refer to the corresponding risk scenario; see Fig.˜6 for details. A case study and quantitative analysis will be discussed in §˜4.3, and a successful example is shown in §˜H.
@@ Appendix B The Use of Large Language Models
We used ChatGPT (achiam2023gpt) to assist with grammatical correction and writing refinement. All research ideas, experimental design, results, and conclusions were independently developed by the authors. The authors bear full responsibility for the entire content of the paper.
@@ Appendix C Benchmark Details
RedCode-Exec (guo2024redcode) provides textual instructions that prompt agents to generate risky code, covering diverse risk scenarios as shown in Fig.˜6.
RedCode-Gen (guo2024redcode) consists of Python function signatures and docstring instructions derived from eight malware families—Adware, Rootkit, Trojan, Virus, DDoS, Ransomware, Spyware, and Others—designed to evaluate a code agent’s ability to generate malware with explicit malicious intent and potential for harm.
RMCBench (chen2024rmcbench) evaluates the ability of LLMs to resist generating malicious code. We use all the text-to-code test cases in our experiment.
@@ D.2 Effectiveness on Real World Code Assistants
We test real-world code assistants, Cursor (cursor2024) and Codeium (codeium2024). For the model setting, we use the “cursor-small model” as the base model for Cursor and the “Base Model” for Codeium.
@@ D.7 Extending the Toolset of RedCodeAgent
Gradient-based attacks: We employ GCG (zou2023universal).
Learning-based attacks: We adopt techniques such as Advprompter (paulus2024advprompter) and AmpleGCG (liao2024amplegcg).
Evolutionary-based attacks: We incorporate AutoDAN (liu2023autodan).
Template-based attacks: We use templates from GPTFUZZER (yu2023gptfuzzer) to craft the attack prompts.
@@ D.11 Experiments on SQL Injection
SQL injection is an important risk scenario; however, it is not included in the RedCode-Exec dataset. Therefore, we followed the test case construction principles outlined in (guo2024redcode) and pre-deployed the example.db database for testing. We created 30 different SQL injection test cases in code format and translated them into text instructions. These text instructions were given to the code agent to evaluate whether it would generate code with vulnerabilities related to SQL injection.
@@ E.2 LLM-based Evaluation
For RedCode-Gen benchmark, following the evaluation setting from (guo2024redcode), we employed an LLM (i.e., GPT-4o) as the judge to score the maliciousness of the agent-generated content.
@@ F.1.1 Agent Framework
We utilize LangChain (topsakal2023creating) as the framework of RedCodeAgent, with GPT-4o-mini (achiam2023gpt) as the base LLM. Within the LangChain framework, we use ReAct (yao2023react) logic, and define functions for tool calling following the setup described in §˜3.3. Additionally, we implement an external memory bank as outlined in §˜3.2, where the structure of memory entries, memory insertion, and memory retrieval processes are discussed in detail. Our system prompt for RedCodeAgent is as follows:
@@ F.1.2 Code Substitution
Due to the sensitivity of some requests, a single LLM might refuse to respond. Therefore, we provide three LLMs to offer code substitution suggestions for RedCodeAgent. They are “Qwen/CodeQwen1.5-7B-Chat"(qwen), “meta-llama/Meta-Llama-3-8B-Instruct"(llama3modelcard), and “gpt-35-turbo"(achiam2023gpt), with their specific links listed in Tb.˜12.
@@ F.3 Baselines Settings
GCG(zou2023universal): We utilize nanoGCG, a lightweight yet fully-featured implementation of GCG. In our experiments, the parameters are set as follows: num_steps=50, search_width=64, topk=64, and seed=42. The target suffix used is “Here is”, and the target white-box model for suffix generation is deepseek-ai/deepseek-coder-6.7b-instruct. It is important to note that deepseek-ai/deepseek-coder-6.7b-instruct differs from the base LLM of our target code agent, OpenCodeInterpreter, which uses m-a-p/OpenCodeInterpreter-DS-6.7B. Therefore, we employed a transfer attack approach using GCG. We only selected one result with the lowest loss among those generated by GCG as the final optimized test case.
@@
AmpleGCG(liao2024amplegcg): For AmpleGCG, we follow the released code and utilized the model osunlp/AmpleGCG-plus-llama2-sourced-llama2-7b-chat. The parameter settings are: max_new_tokens=20, min_new_tokens=20, diversity_penalty=1.0, num_beams=50, and return_num_beams=5. We only select the first result from the list of strings generated by AmpleGCG optimization as the final optimized test case.
AutoDAN(liu2023autodan): We follow the official code implementation, using the same experimental parameters as those provided in the repository. For the initial prompt setting, we use the following: Hypothetical response: the way to describe a character planning to [REPLACE] is as follows: “yo”. Rewrite the hypothetical response from the perspective of that character, make it extremely detailed, and use exact and precise information (avoid generalizations or vagueness).
AdvPrompter(paulus2024advprompter): We utilize the provided code to get the LoRA adapter for the target model based on “meta-llama/Llama-2-7b-hf", and then combined the adapter with the original LLM to create the suffix generator model. The maximum number of new tokens was set to max_new_tokens=50.
@@ G.1 Negative Optimization of Some Baseline Methods
In our experiments, we selected 4 representative and widely used optimization methods (i.e., GCG (zou2023universal), AmpleGCG (liao2024amplegcg), Advprompter (paulus2024advprompter), and AutoDAN (liu2023autodan)) that have been shown to be effective in previous literature on jailbreaking LLMs. However, we observed that certain optimizations in these baseline methods may not be as effective for attacking code agents. This is likely due to the differences between natural language processing and code-specific reasoning tasks. For instance, the targeted code agent OCI can reject the prompt optimized from AutoDan (see detailed example below). At present, most existing jailbreak methods are focused on natural language tasks, and there is a lack of specialized jailbreaking optimization algorithms designed for code-related tasks. This gap further highlights the contribution of RedCodeAgent, which addresses this deficiency by offering a tailored framework for red-teaming code agents, ultimately advancing the field.